# Direct Unlearning Optimization for Robust and Safe Text-to-Image Models

### v1

@@ Abstract
Recent advancements in text-to-image (T2I) models have greatly benefited from large-scale datasets, but they also pose significant risks due to the potential generation of unsafe content. To mitigate this issue, researchers have developed unlearning techniques to remove the model’s ability to generate potentially harmful content. However, these methods are easily bypassed by adversarial attacks, making them unreliable for ensuring the safety of generated images. In this paper, we propose Direct Unlearning Optimization (DUO), a novel framework for removing Not Safe For Work (NSFW) content from T2I models while preserving their performance on unrelated topics. DUO employs a preference optimization approach using curated paired image data, ensuring that the model learns to remove unsafe visual concepts while retaining unrelated features. Furthermore, we introduce an output-preserving regularization term to maintain the model’s generative capabilities on safe content. Extensive experiments demonstrate that DUO can robustly defend against various state-of-the-art red teaming methods without significant performance degradation on unrelated topics, as measured by FID and CLIP scores. Our work contributes to the development of safer and more reliable T2I models, paving the way for their responsible deployment in both closed-source and open-source scenarios.
@@ 1 Introduction
In recent years, text-to-image (T2I) models [21, 46, 44, 39, 8, 7] have experienced significant advancements thanks to large-scale data and diffusion models. However, the large-scale web-crawled data, e.g., LAION [50] often include a significant amount of inappropriate and objectionable material, so T2I models trained on such data may pose a potential risk of generating harmful content such as Not Safe For Work (NSFW) content, copyright infringement, and a violation of privacy [45, 4]. Filtering and curating such large-scale training data [54] is not feasible for large-scale datasets. At the production level, service providers usually block inappropriately generated images with ad-hoc classifier [54], but attackers can easily bypass the classifier by accessing the model directly.
@@
To address the generation of unsafe content, there has been a substantial amount of research focusing on directly fine-tuning diffusion models to forget unsafe concepts [11, 12, 25, 32, 16, 35, 62]. Recent work in the field has primarily focused on prompt-based approaches to unlearning, which aim to induce models to behave as if they had been provided with safe prompts when an unsafe prompt is given [11, 25, 32, 62, 35, 12]. These methods effectively censor NSFW image generation without degrading the image generation capability. However, recent studies [41, 55, 59] warn that by introducing slight perturbations to the unsafe prompts, it is possible to circumvent these censoring mechanisms and generate undesirable images (Figure 1).
@@
Figure 1: Image-based unlearning. Prompt-based unlearning can be easily circumvented with adversarial prompt attack. On the other hand, image-based unlearning robustly produces safe images regardless of the given prompt. We use Refer to caption for publication purposes.
@@
Current unlearning methods are vulnerable to adversarial attacks because it is difficult to cover all NSFW concepts using prompts alone. For instance, even if the prompt “nudity” is removed, many related concepts like “naked,” “erotic,” or “sexual” may remain. Furthermore, it is infeasible to anticipate and eliminate all possible adversarial prompts beforehand. Moreover, unlearning methods that rely on text conditioning cause the model to behave as if it cannot comprehend the unsafe prompts, rather than removing the internal visual concepts within the model.
@@
To fundamentally address limitations, it is necessary to guide the model to remove unsafe visual features regardless of the prompt. One way to achieve this is by directly unlearning the unsafe images instead of prompts. However, simply moving away from the target image leads to ambiguity regarding which concepts to forget and can cause forgetting of unrelated visual features. For instance, if we make the model unlearn the image in Figure 3-(a), it would lose the ability to generate not only nudity but also unrelated concepts like women or forests.
@@
In this paper, we propose Direct Unlearning Optimization (DUO), a method to remove unsafe visual features directly from the model while maintaining generation quality for unrelated topics. Specifically, we created paired data by using unsafe images and their counterparts, from which unsafe concepts were removed using the SDEdit [33]. Then, we employ Direct Preference Optimization (DPO) [43] to make the model favor the latter group while encouraging it to move away from the former group. As shown in Figure 3-(b), preference optimization through a paired dataset explicitly specifies which features of a given image should be forgotten. To further regularize the model not to unlearn for unrelated topics, we introduce output-preserving regularization [36, 61, 6] that forces to preserve the denoising capability of the model. We call our framework as Direct Unlearning Optimization (DUO). We demonstrate that our method can robustly defend against various state-of-the-art red teaming methods [55, 41] without significant performance degradation in unrelated topics, as measured by LPIPS [64], FID [18] and CLIP scores [42]. Additionally, through an ablation study, we validate the efficacy of output preservation regularization and visual feature unlearning.
@@ 2 Related work
The safety mechanisms designed to prevent undesirable behavior in T2I models can be classified into four categories: dataset filtering [54, 14], concept unlearning [11, 12, 25, 32, 16, 35, 62], in-generation guidance [49], and post-generation content screening [1, 54, 3]. These methods are orthogonal to each other, and using them together can make the T2I generation system safer.
Due to the advantage of not requiring training from scratch, fine-tuning-based unlearning has been actively researched recently. Pioneering works like ESD [11] and CA [25] train the model to generate the same noise regardless of whether an unsafe prompt is present or not. SPM [32] trains a one-dimensional regulator instead of full model training. TIME [35], UCE [12] and MACE [30] update prompt-dependent cross-attention model weights with closed-form equation. Forget-me-not [62] suppresses the attention map of cross-attention layers for unsafe prompts. SA [16] proposed an unlearning method based on continual learning, utilizing Elastic Weight Consolidation (EWC). These works rely on prompts during training, making them vulnerable to adversarial prompt attacks, as discussed later.
Meanwhile, SafeGen [28], as a concurrent work with ours, addresses similar concerns regarding adversarial prompts and utilizing only images for training. They use blurred images and guide a model using supervised learning. In contrast to their work, we utilize paired data generated with SDEdit [33] to enable the model to erase only unsafe concepts and employ preference optimization. Therefore, we have the advantage of not requiring separate benign data and producing erased results more closely resemble real images.
@@
Red teaming is a method for searching for vulnerabilities in security systems. Extending this concept to machine learning, many red teaming techniques have been proposed to explore how robust a model’s safety mechanism is. For Text-to-Image (T2I) models, studies have shown that even with safety mechanisms applied, it is still possible to generate harmful images through prompt engineering [55, 41, 59, 9, 66, 15]. Ring-A-bell [55] and SneakyPrompt [59] generate adversarial prompts in a black box scenario without access to the denoising model. They collect unsafe prompts and find prompts with similar embeddings through optimization. On the other hand, Concept Inversion [41] is a method that can be used in a white box scenario where access to the gradient of the denoising model is available. This method learns a special token $\langle c\rangle$ through textual inversion that contains information about unsafe images, and then uses it to generate unsafe images. The goal of our research is to propose a safety mechanism that is robust against red teaming.
@@ 3.2 Synthesizing paired image data to resolve ambiguity in image-based unlearning
Figure 5: Effectiveness of utilizing SDEdit for generating paired image data for unlearning. When unlearning (a), if a retain sample is naively generated as shown in (b), the model cannot accurately determine what information should be retained and what should be forgotten. Using SDEdit as shown in (c), the model can identify what information should be retained and what should be forgotten from the undesirable sample. We use Refer to caption for publication purposes.
As we mentioned in Section 1, even images with unsafe concepts contain numerous unrelated visual features that need to be preserved (e.g., photo-realism, background, and humans). Therefore, a method is needed to selectively remove only the visual features associated with the targeted unsafe concept. Intuitively, if we provide the model with images that have identical unrelated visual features but differ in the targeted unsafe concept, the model can be guided to unlearn only the features related to the unsafe concept. Unfortunately, as shown in Figure 5 (b), naively generating paired images with different prompts often resulted in changes to the unrelated attributes of the images.
@@ 3.3 Concept unlearning as a preference optimization problem
where $p_{\phi}$ denotes the pretrained model distribution, and the hyperparameter $\beta$ controls the extent to which the model diverges from the prior distribution. The above equation has a unique closed-form solution $p_{\theta}^{*}$ [13, 24, 38, 40]:
@@
By leveraging Jensen’s inequality and the ELBO, this loss can be expressed as a combination of fully tractable denoising score matching losses (see the Appendix C for detailed derivation):
@@ 3.4 Output-preserving regularization
Unfortunately, in preliminary experiments, we discovered that DPO’s KL divergence regularization alone is insufficient for prior preservation. The main reason was that the gradient ascent term in Eq. (17) does not simply unlearn $x_{0}^{+}$, but also decreases the model’s denoising ability.
To mitigate this issue, we introduced the following novel regularization term to maintain the diffusion model’s denoising capability:
@@
The above regularization maintains the output even when unlearning is performed and the image is completely noise, i.e., at $x_{T}$. The reason for limiting it to $t=T$ is that when image information remains, output preservation can interfere with unlearning. On the other hand, when $t=T$, only complete noise information remains in the image, so we expected that enforcing output preservation would not affect unlearning.
@@ 4.1 Experiments setup
We use Stable Diffuson 1.4v (SD1.4v) with a LoRA [22, 47] with the Adam optimizer for fine-tuning. [11] We do not train the cross-attention layers to minimize the dependency on the text prompt for unlearning. Our experiments are conducted on Stable Diffusion version 1.4 to compare with other baseline methods. We use ESD [11], UCE [12], SPM [32] and MACE [30] as baselines for the concept unlearning category of the safety mechanism. SPM has a Facilitated Transfer (FT) mechanism that adjusts the strength of unlearning based on the similarity of given prompt to the unlearned concepts. This mechanism can be disadvantageous against adversarial prompt attacks, so we additionally report results from SPM without this mechanism (SPM w/o FT). For implementation details, please refer to the Appendix.
@@
To classify whether images contain violence, we used GPT-4o, a state-of-the-art multimodal LLM [57, 60, 1]. We use the user prompt, "Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? Please answer yes or no." [48] along with the image. If the model’s response is “yes”, we determine that it contains violence, while “no” indicated unlearning success. If neither token is present, we regenerate the model output. For more discussion on violence evaluation, please refer to the appendix (WIP).
To evaluate model performance unrelated to unlearned concept, we measure the FID [18] and CLIP scores [17] using MS COCO 30k validation dataset [29]. These metric provide insights into the capabilities of unlearned models. Additionally, we report the LPIPS [64] between images from the original SD1.4v model with those from the unlearned model using identical noise. This quantifies how the unlearned model’s distribution differs from the original.
@@ 4.2 Red teaming results
Figure 8: Qualitative result on nudity. Ours We use Refer to caption for publication purposes.
Unlearning is a task to satisfy the trade-off between unlearning performance and prior preservation simultaneously. Therefore, we report the Pareto curve according to the KL regularization. We plot the Pareto curve by adjusting the magnitude of KL regularization. Specifically, each point in the figures corresponds to $\beta\in\{500,250,100,50\}$.
Nudity.
Figure 6 shows the quantitative results for nudity detection. Our method maintains model performance comparable to the current state-of-the-art SPM in terms of prior preservation while achieving a defense success rate (DSR) of nearly 90% against all red teaming methods. A notable characteristic of the Pareto curve is the sharp bend where the increase in unlearning performance diminishes and the model rapidly loses its prior knowledge. Based on this observation, we recommend choosing $\beta=100$.
Figure 8 presents the qualitative results from Ring-A-Bell and concept inversion attacks. While previous prompt-based unlearning methods are easily circumvented with red teaming, our method demonstrates robust defense against these attacks. This result highlights the effectiveness of our image-based unlearning approach, which is independent of the prompt and robust against prompt-based red teaming. Interestingly, in the case of concept inversion, completely corrupted images are generated, unlike in the Ring-A-Bell attack. Through experimentation, we observed that it is challenging to robustly block textual inversion unless the generated result deviates significantly from the data manifold. This suggests that to defend against white box red teaming, which leverages gradient information, the internal features of the model must be completely removed rather than merely bypassed. Analyzing the different mechanisms of black box and white box red teaming presents an interesting direction for future research.
In Table 2, we report the FID and CLIP scores for MS COCO 30k. These scores indicate how well the unlearned model preserves its ability to generate images for unrelated prompts. Unlike other methods, DUO does not modify text embeddings or prompt-dependent weights; instead, it removes visual features directly from the model, making prior preservation quite challenging. Nonetheless, in terms of FID and CLIP scores, DUO shows performance comparable to other unlearning methods. Thus, for nudity, DUO is robust against prompt-based adversarial attacks while also preserving the original model’s generation capabilities for unrelated concepts. Figure 9 qualitatively confirms that the generation results do not significantly differ between the prior model and the model after unlearning.
@@
ESD 14.05 29.97
UCE 14.30 30.64
SPM 13.74 30.92
SPM [^2] SPM without facilitate transfer 14.15 30.80
MACE 13.51 29.06
DUO 14.16 30.84
@@
ESD 21.33 27.35
UCE 16.20 30.40
SPM 13.60 30.89
SPM 13.44 30.84
MACE 14.85 29.14
DUO 17.52 30.34
Figure 10: Quantitative result on violence. The defense success rate (DSR) refers to the proportion of desirable concepts are generated. Prior preservation represents 1 - LPIPS between images generated by the prior model and the unlearned model. Results closer to the top right indicate better outcomes.
@@
Figure 10 demonstrates the red teaming result on violence. Similar to nudity, DUO shows remarkable robustness against attacks while maintaining the same prior preservation score. We observe that the defense success rate values are generally not completely blocked compared to nudity. We believe is due to the fact that violence is a more abstract and complex concept compared to nudity, making unlearning more difficult. Figure 12 show qualitative results. We present the outcomes using $\beta=100$. Table 2 shows the FID and CLIP scores for MS COCO 30k. Similar to the nudity concept removal experiment, our method maintains comparable prior preservation performance to existing methods. Ulike the nudity experiment, in the violence concept removal experiment, we observe that the trade-off between prior preservation and defense success rate varies more steeply with the parameter $\beta$.

### v2

@@ Abstract
Recent advancements in text-to-image (T2I) models have unlocked a wide range of applications but also present significant risks, particularly in their potential to generate unsafe content. To mitigate this issue, researchers have developed unlearning techniques to remove the model’s ability to generate potentially harmful content. However, these methods are easily bypassed by adversarial attacks, making them unreliable for ensuring the safety of generated images. In this paper, we propose Direct Unlearning Optimization (DUO), a novel framework for removing Not Safe For Work (NSFW) content from T2I models while preserving their performance on unrelated topics. DUO employs a preference optimization approach using curated paired image data, ensuring that the model learns to remove unsafe visual concepts while retaining unrelated features. Furthermore, we introduce an output-preserving regularization term to maintain the model’s generative capabilities on safe content. Extensive experiments demonstrate that DUO can robustly defend against various state-of-the-art red teaming methods without significant performance degradation on unrelated topics, as measured by FID and CLIP scores. Our work contributes to the development of safer and more reliable T2I models, paving the way for their responsible deployment in both closed-source and open-source scenarios.
@@ 1 Introduction
In recent years, text-to-image (T2I) models [22, 47, 45, 40, 8, 7] have experienced significant advancements thanks to large-scale data and diffusion models. However, the large-scale web-crawled data such as LAION [51] often include a significant amount of inappropriate and objectionable material, so T2I models trained on such data may pose a potential risk of generating harmful content including Not Safe For Work (NSFW) content, copyright infringement, and a violation of privacy [46, 4]. Filtering and curating such large-scale training data [55] is not feasible for large-scale datasets. At the production level, service providers usually block inappropriately generated images with an ad-hoc classifier [55], but the classifier can block normal images due to false positive detections and can be easily bypassed if the model weights are open-sourced.
@@
To address the generation of unsafe content, there has been a substantial amount of research focusing on directly fine-tuning diffusion models to forget unsafe concepts [12, 13, 26, 33, 17, 36, 63]. Recent work in the field has focused mainly on prompt-based approaches to unlearning, which aim to induce models to behave as if they had received safe prompts when an unsafe prompt is given [12, 26, 33, 63, 36, 13]. These methods effectively censor NSFW image generation without degrading the image generation capability. However, recent studies [42, 56, 60] warn that by introducing slight perturbations to the unsafe prompts using adversarial prompts [56, 60] or embeddings[42], it is possible to circumvent these censoring mechanisms and generate undesirable images (Figure 1).
@@
Figure 1: Visualization of the advantages of image-based unlearning. Prompt-based unlearning can be easily circumvented with adversarial prompt attack. On the other hand, image-based unlearning robustly produces safe images regardless of the given prompt. We use Refer to caption for publication purposes.
@@
Current unlearning methods are vulnerable to adversarial attacks because there are many synonyms or indirect expressions for NSFW concepts in text. For instance, even if the prompt “nudity” is removed, many related concepts like “naked,” “erotic,” or “sexual” may remain. Furthermore, it is infeasible to anticipate and eliminate all possible adversarial prompts beforehand. Moreover, unlearning methods that rely on text conditioning cause the model to behave as if it cannot comprehend the unsafe prompts, rather than removing the internal visual concepts within the model.
@@
To fundamentally address limitations, it is necessary to guide the model to guide the model to remove the image-related features that produce unsafe images, regardless of the prompt. One way to achieve this is by directly unlearning the unsafe images instead of prompts. However, simply moving away from the target image leads to ambiguity regarding which concepts to forget and can cause forgetting of unrelated visual features. For instance, if we make the model unlearn the image in Figure 3-(a), it would lose the ability to generate not only nudity but also unrelated concepts like women or forests.
@@
In this paper, we propose Direct Unlearning Optimization (DUO), a method to remove unsafe visual features directly from the model while maintaining generation quality for unrelated topics. Specifically, we created paired data by using unsafe images and their counterparts, from which unsafe concepts were removed using SDEdit [34] (Section 3.2). Then, we employ Direct Preference Optimization (DPO) [44] to make the model favor the latter group while encouraging it to move away from the former group (Section 3.3). As shown in Figure 3-(b), preference optimization through the paired dataset explicitly specifies which features of a given image should be forgotten. To further regularize the model not to unlearn for unrelated topics, we introduce output-preserving regularization [37, 62, 6] that forces to preserve the denoising capability of the model. We call our framework as Direct Unlearning Optimization (DUO). We demonstrate that our method can robustly defend against various state-of-the-art red teaming methods [56, 42] without significant performance degradation in unrelated topics, as measured by LPIPS [65], FID [19] and CLIP scores [43]. Additionally, through an ablation study, we validate the efficacy of output preservation regularization and visual feature unlearning.
@@ 2 Related work
The safety mechanisms designed to prevent undesirable behavior in T2I models can be classified into four categories: dataset filtering [55, 15], concept unlearning [12, 13, 26, 33, 17, 36, 63], in-generation guidance [50], and post-generation content screening [1, 55, 3]. These methods are orthogonal to each other and have different limitations. In this work, we focus on concept unlearning and can combine it with other techniques to ensure more secure T2I generation.
Due to the advantage of not requiring training from scratch, fine-tuning-based unlearning has been actively researched recently. Pioneering works like ESD [12] and CA [26] train the model to generate the same noise regardless of whether an unsafe prompt is present or not. SPM [33] trains one-dimensional adapters that regularize the model instead of full model training. TIME [36], UCE [13] and MACE [31] update the prompt-dependent cross-attention weights with a closed-form equation. Forget-me-not [63] suppresses the attention map of cross-attention layers for unsafe prompts. SA [17] proposed an unlearning method based on continual learning, utilizing Elastic Weight Consolidation (EWC). These works rely on prompts during training, making them vulnerable to adversarial prompt attacks, as discussed later.
Meanwhile, SafeGen [29], as a concurrent work with ours, addresses similar concerns regarding adversarial prompts and utilizing image pairs to remove unsafe features for training. They use blurred unsafe images to guide the model using supervised learning, ensuring that unsafe images are generated in a blurred form. In contrast to their work, we utilize paired data generated with SDEdit [34] to enable the model to erase only unsafe concepts and employ preference optimization. Therefore, our method can provide more selective guidance for the visual features that need to be removed, and avoid direct training on blurred images, which can potentially harm the generation capabilities.
@@
Red teaming is a method for searching for vulnerabilities in security systems. Extending this concept to machine learning, many red teaming techniques have been proposed to explore how robust a model’s safety mechanism is. For Text-to-Image (T2I) models, studies have shown that even with safety mechanisms applied, it is still possible to generate harmful images through prompt engineering [56, 42, 60, 9, 67, 16]. Ring-A-bell [56] and SneakyPrompt [60] generate adversarial prompts in a black-box scenario without access to the diffusion model. They collect unsafe prompts and find prompts with similar embeddings through optimization. On the other hand, Concept Inversion [42] is a method that can be used in a white box scenario where access to the gradient of the diffusion model is available. This method learns a special token $\langle c\rangle$ through textual inversion [42] that contains information about unsafe images, and then uses it to generate unsafe images. The goal of our research is to propose a safety mechanism that is robust against red teaming.
@@ 3.2 Synthesizing paired image data to resolve ambiguity in image-based unlearning
Figure 5: Effectiveness of utilizing SDEdit for generating paired image data for unlearning. When unlearning unsafe images (a), we use safe images (b, c) to indicate which visual features should be retained. While prompt substitution (b) prevents the model from accurately determining what visual features to retain or forget, SDEdit (c) enables the model to identify which information from the undesirable sample should be kept or discarded. We use Refer to caption for publication purposes.
As we mentioned in Section 1, even images with unsafe concepts contain numerous unrelated visual features that need to be preserved (e.g., photo-realism, background, and humans). Therefore, a method is needed to selectively remove only the visual features associated with the targeted unsafe concept. Intuitively, if we provide the model with images that have identical, unrelated visual features but differ in the targeted unsafe concept, the model can be guided to unlearn only the features related to the unsafe concept. Unfortunately, as shown in Figure 5 (b), naively generating paired images by replacing unsafe words with safe words in prompts often resulted in changes to the unrelated attributes of the images.
@@ 3.3 Concept unlearning as a preference optimization problem
where $p_{\phi}$ denotes the pretrained model distribution, and the hyperparameter $\beta$ controls the extent to which the model diverges from the prior distribution. The above equation has a unique closed-form solution $p_{\theta}^{*}$ [14, 25, 39, 41] (see Appendix D.1 for detailed derivation):
@@
By leveraging Jensen’s inequality and the ELBO, this loss can be expressed as a combination of fully tractable denoising score matching losses (see Appendix D.2 for detailed derivation):
@@ 3.4 Output-preserving regularization
Unfortunately, in preliminary experiments, we discovered that DPO’s KL divergence regularization alone is insufficient for prior preservation. To mitigate this issue, we introduced the following novel regularization term to maintain the diffusion model’s denoising capability:
@@
The above regularization maintains the output even when unlearning is performed, and the image is completely noise, i.e., at $x_{T}$. The reason for limiting it to $t=T$ is to prevent output preservation regularization from interfering with the removal of the knowledge about unsafe visual features contained in the image $x_{0}$.
@@ 4.1 Experiments setup
We use Stable Diffuson 1.4v (SD1.4v) with a LoRA [23, 48] rank of 32 with the Adam optimizer for fine-tuning. We observe that it is difficult to unlearn violence with a single LoRA, because it has various subcategories. To address this, we decompose it into four distinctive concepts: blood, suffering, gun, and horror. We apply DUO to each concept and merge them to use as violence unlearning LoRA. We do not train the cross-attention layers to minimize the dependency on the text prompt for unlearning [12]. Our experiments are conducted on Stable Diffusion version 1.4 to compare with other baseline methods. We use ESD [12], UCE [13], SPM [33] as baselines for the concept unlearning category of the safety mechanism. For implementation details, please refer to the Appendix.
@@
To classify whether images contain violence, we used GPT-4o, a state-of-the-art multimodal LLM [58, 61, 1]. We use the user prompt, "Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? Please answer yes or no." [49] along with the image. If the model’s response is “yes”, we determine that it contains violence, while “no” indicated unlearning success. If neither token is present, we regenerate the model output.
To evaluate model performance unrelated to the unlearned concept, we measure the FID [19] and CLIP scores [18] using MS COCO 30k validation dataset [30]. These metrics provide insights into the capabilities of unlearned models. Additionally, we report the LPIPS [65] between the images from the original SD1.4v model and those from the unlearned model using identical noises. This quantifies how the unlearned model’s distribution differs from the original.
@@ 4.2 Red teaming results
Figure 8: Qualitative result on nudity. We used $\beta=500$ for Ring-A-bell and $\beta=250$ for Concept Inversion. We use Refer to caption for publication purposes.
Unlearning is a task to satisfy the trade-off between unlearning performance and prior preservation simultaneously. We plot the Pareto curve by adjusting the magnitude of KL regularization. Specifically, each point in the figures corresponds to $\beta\in\{100,250,500,1000,2000\}$.
@@
SD (1.4v) 13.52 30.95
ESD 14.07 30.00
UCE 13.95 30.89
SPM 14.05 30.84
DUO ($\beta=500$) 13.65 19.88
DUO ($\beta=250$) 13.59 30.84
@@
SD (1.4v) 13.52 30.95
ESD 16.87 29.44
UCE 14.04 30.74
SPM 13.53 30.93
DUO ($\beta=1000$) 13.37 30.78
DUO ($\beta=500$) 18.28 30.18
Nudity.
Figure 6 shows the quantitative results for nudity detection. Our method maintains model performance comparable to the current state-of-the-art SPM in terms of prior preservation while achieving a defense success rate (DSR) of nearly 90% against all red teaming methods. A notable characteristic of the Pareto curve is the sharp bend where the increase in unlearning performance diminishes, and the model rapidly loses its prior knowledge. Based on this observation, we recommend choosing $\beta=500$ and $\beta=250$ for defending against black-box and white-box red teaming methods, respectively.
Figure 8 presents the qualitative results from Ring-A-Bell and concept inversion attacks. While previous prompt-based unlearning methods are easily circumvented with red teaming, our method demonstrates robust defense against these attacks. This result highlights the effectiveness of our image-based unlearning approach, which is independent of the prompt and robust against prompt-based red teaming. Interestingly, in the case of concept inversion, completely corrupted images are generated, unlike in the Ring-A-Bell attack. Through experimentation, we observed that it is challenging to robustly block textual inversion unless the generated result deviates significantly from the data manifold. This suggests that to defend against white box red teaming, which leverages gradient information, the internal features of the model must be completely removed rather than merely bypassed. Analyzing the different mechanisms of black-box and white-box red teaming presents an interesting direction for future research.
In Table 2, we report the FID and CLIP scores for MS COCO 30k. These scores indicate how well the unlearned model preserves its ability to generate images for unrelated prompts. Unlike the other methods, DUO does not modify text embeddings or prompt-dependent weights; instead, it removes visual features directly from the model, making prior preservation quite challenging. Nonetheless, regarding FID and CLIP scores, DUO shows performance comparable to other unlearning methods. Thus, for nudity, DUO is robust against prompt-based adversarial attacks while also preserving the original model’s generation capabilities for unrelated concepts. Figure 9 qualitatively confirms that the generation results do not significantly differ between the prior model and the model after unlearning.
Figure 10: Quantitative result on violence. The defense success rate (DSR) refers to the proportion of desirable concepts generated. Prior preservation represents 1 - LPIPS between images generated by the prior model and the unlearned model. Results closer to the top right indicate better outcomes.
@@
Figure 10 demonstrates the red teaming result on violence. Like nudity, DUO shows remarkable robustness against attacks while maintaining the same prior preservation score. We observe that the defense success rate values are generally not completely blocked compared to nudity. We believe this is due to the fact that violence is a more abstract and complex concept compared to nudity, making unlearning more difficult. Figure 12 show qualitative results. We present the outcomes using $\beta=500$ for ring-a-bell and $\beta=1000$ for concept inversion. Table 2 shows the FID and CLIP scores for MS COCO 30k. Similar to the nudity concept removal experiment, our method maintains comparable prior preservation performance to existing methods. Unlike the nudity experiment, in the violence concept removal experiment, we observe that the trade-off between prior preservation and defense success rate varies more steeply with the parameter $\beta$.